OpenStudy (rissa):

Hi is there anyone who can explain the steps to solve radical with index of 3 8x^2/ radical x

OpenStudy (evoker):

It is this \[\sqrt[3]{8x^2}/\sqrt[3]{x}\]

OpenStudy (rissa):

yes

OpenStudy (evoker):

by solve I assume they mean simplify

OpenStudy (evoker):

so anyways since they have the same radical power first they can be combined to give \[\sqrt[3]{8x}\]

OpenStudy (rissa):

yeah i know i multiply by the conjugate but then i a little lost

OpenStudy (evoker):

ok now cube root means you are looking for triple copies inside the radical, 8=2*2*2 so you can pull out a factor of 2

OpenStudy (evoker):

So simplified you would have \[2\sqrt[3]{x}\]

OpenStudy (rissa):

im confused on how you got rid of the \[x ^{2}\]

OpenStudy (evoker):

\[\sqrt[3]{8x^2}/\sqrt[3]{x}=\sqrt[3]{\frac{ 8x^2 }{ x }}\]